# Transverse Phase Space Tomography at FACET-II

### v1

@@ 3 Two bunch Configuration
We further mapped out the phase space with roughly 0.5 nC, 1 nC, 1.5 nC, and 2 nC bunch charge, respectively, and the result is shown in figure 3. We observe that the higher the bunch charge, the more stretched out the phase space is spatially.
In the horizontal plane, the result does not agree with our preliminary simulation result. We attributed the mismatch to three effects and their interplay: (1) transverse space-charge force; (2) solenoid setting for emittance compensation; and (3) laser interference. During the experiment, the solenoid setting is unchanged, so it is not efficiently compensating bunches with different charges. The laser pulse time separation is zero, but it is split and recombined, thus the laser recombination effect may also play a role. A comprehensive simulation and analysis are in progress to understand the measurement result.

### v2

@@ 3 Two bunch Configuration
We further mapped out the phase space with roughly 0.5 nC, 1 nC, 1.5 nC, and 2 nC bunch charge, respectively, and the result is shown in figure 3. We observe that the higher the bunch charge, the more spatially stretched the phase space is.
In the horizontal plane, the result does not agree with our preliminary simulation result. We attributed the mismatch to three effects and their interplay: (1) space-charge force; (2) solenoid setting for emittance compensation; and (3) collective effects in the laser heater chicane. During the experiment, the solenoid setting is unchanged, so it is not efficiently compensating bunches with different charges. Collective effects in the laser heater chicane can play a major role, since the horizontal phase space is particularly affected. A comprehensive simulation and analysis are in progress to understand the measurement result.